OpenStudy (princeevee):

x/6 divided by 1/6 A. 6x B. x C. 1 D. 6

Nnesha (nnesha):

change division to multiplication multiply the top by reciprocal of the bottom fraction \[\huge\rm \frac{ \frac{ x }{ 6 } }{ \frac{ 1 }{ 6 } }\] what is reciprocal of 1/6 ?

OpenStudy (princeevee):

?

OpenStudy (princeevee):

6?

Nnesha (nnesha):

here is an example \[\frac{ \frac{ 2 }{ 3 } }{ \frac{ a }{ b } } = \frac{ 2 }{ 3 } \times \frac{ b }{ a }\]

Nnesha (nnesha):

yes right so multiply top fraction x/6 by 6\[\frac{ x }{ 6 } \times 6 =??\]

Nnesha (nnesha):

multiply the numerator by numerator and denominator by denominator 6 is same as 6/1 \[\frac{ x }{ 6 } \times \frac{6}{1} =?\]

OpenStudy (princeevee):

6X?

Nnesha (nnesha):

hmm  now divide 6x/6 =?